a grocery store bought milk for $2.60 per half gallon and stored it in two refrigerators. During the night, one refrigerator malfunctioned and ruined 12 half gallons. if the remaining milk is sold for $4.02 per half gallon, how many half gallons did the store buy if they made a profit of $88.08
Answer by lwsshak3(11628)

a grocery store bought milk for $2.60 per half gallon and stored it in two refrigerators. During the night, one refrigerator malfunctioned and ruined 12 half gallons. if the remaining milk is sold for $4.02 per half gallon, how many half gallons did the store buy if they made a profit of $88.08
**
let x=number of half gallons of milk the store bought
x-12=remaining half gallons
profit=revenue-cost
88.08=4.02(x-12)-2.60x
88.08=4.02x-48.24-2.6x
1.42x=88.08+48.24=136.32
x=136.32/1.42=96
ans:
number of half gallons of milk the store bought=96